OpenStudy (anonymous):

Solving Radical Equations -4 √x+9(end of SQR)=20

OpenStudy (bahrom7893):

square both sides

OpenStudy (bahrom7893):

16(x+9)=400

OpenStudy (bahrom7893):

16x+9*16=400 16x = 400-9*16 x = (400-9*16)/16

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

The answer is wrong.